# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v2

@@ 4.7.2 Summary of UC Security
  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$
  • Profile P2 (MPC): UC security against dishonest majority via MPC (Theorem 6.2). This provides the strongest cryptographic guarantees with no hardware trust.
  • Profile P3 (2PC): UC security under 1-of-2 CP honest assumption (Theorem 6.5). Combines cryptographic security with best performance.
The choice of profile affects trust assumptions and performance, but not the fundamental cryptographic security of the resulting signatures.
@@ 5.1 Implementation Overview
  • Polynomial Arithmetic: Operations over $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$ with $q=8380417$, including addition, subtraction, multiplication, and NTT-based fast multiplication.
  • Shamir Secret Sharing: Polynomial evaluation and Lagrange interpolation over $\mathbb{Z}_{q}$. Lagrange coefficients $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$ depend on the signing set $S$ and are computed per-session in $O(T^{2})$ time. For fixed-$S$ deployments (e.g., always using the same $T$ parties), coefficients can be precomputed and cached.
  • Mask Generation: Pairwise-canceling mask computation using SHAKE-256 as the PRF, with efficient seed management.
  • Protocol Logic: Three-round signing protocol with commitment verification, challenge derivation, and signature aggregation.
@@
  1. 1.
    NTT Acceleration: Polynomial multiplication uses the Number Theoretic Transform with $\omega=1753$ as a primitive $256$-th root of unity modulo $q$. This reduces multiplication complexity from $O(n^{2})$ to $O(n\log n)$, providing a $7\text{--}10\times$ speedup.
  2. 2.
    Montgomery Reduction: Modular arithmetic uses Montgomery form for efficient reduction without division.
  3. 3.
    Vectorized Operations: Polynomial and vector operations leverage NumPy’s SIMD-optimized routines.
  4. 4.
    Lazy Reduction: Intermediate polynomial coefficients are reduced modulo $q$ only when necessary to prevent overflow.
@@ 5.2 Benchmark Methodology
All benchmarks were conducted on the following system:
  • CPU: Intel Core i7-12700H (14 cores, 20 threads)
  • RAM: 32 GB DDR5-4800
  • OS: Ubuntu 22.04 LTS
  • Python: 3.11.5 with NumPy 1.26.2
@@
Select configurations were verified with 1000 trials (marked in tables); others use 100 trials. We report median values with 95% confidence intervals.
@@ 5.3 Performance Results
Table 4 shows signing performance for various threshold configurations with $N\approx 2T$.
$T$ NN Time Att. Succ. $1/p_{\text{naive}}$ Speedup
3 5 0.97 s 3.2 31% $10^{2}$ $38\times$
5 9 1.77 s 3.9 26% $10^{3}$ $800\times$
8 15 2.41 s 3.5 28% $10^{5}$ $10^{5}$
12 23 3.15 s 3.2 31% $10^{8}$ $10^{7}$
16 31 4.01 s 3.1 32% $^{11} 1010$$^{10}$
24 47 8.47 s 4.3 23% $10^{16}$ $10^{16}$
32 63 10.5 s 4.0 25% $10^{22}$ $10^{21}$
Table 4: Signing performance (ML-DSA-65, Python, Intel i7). 1000 trials. Success rates remain constant ($\sim$23–32%) vs. naive $(0.2)^{T}$ degradation. Speedup $(1/p_{\text{naive}})/\text{Attempts}$.
Signing time scales linearly with $|S|$, while success rate remains constant at $23\text{--}32\%$ (matching single-signer ML-DSA). Typical configurations complete in 1–4 seconds; production Rust/C implementations would be faster. Round 3 dominates timing ($\approx 58\%$) due to mask computation.
@@ 5.4 Protocol Profile Comparison
Table 5 compares the three deployment profiles with experimental measurements for ML-DSA-44, $(n=5,t=3)$.
Profile Rounds Trust Sign Time Attempts 95% CI
[…]
Table 5: Protocol profile comparison (ML-DSA-44, $n=5$, $t=3$, 50 trials). P3-2PC achieves the best performance with the weakest trust assumption among practical options. Sign time includes all attempts until success.
Profile P3 (2PC-assisted) achieves the fastest signing times while requiring only that one of two computation parties is honest, making it suitable for deployments where TEE hardware is unavailable.
@@ 5.5 Comparison with Prior Work
Scheme $T$ Size Rnd FIPS Time
dPN25 [dPN25] 8 3.3 KB 3 Yes
Noise flood [BCDG21] $\infty$ 17 KB 3 No 0.3 s
Ringtail [BMMSZ25] $\infty$ 13 KB 2 No 0.2 s
Ours (P3) $\infty$ 3.3 KB 3–5 Yes 0.25 s
Ours (P1) $\infty$ 3.3 KB 3 Yes 0.33 s
Ours (P2) $\infty$ 3.3 KB 8 Yes 0.29 s
Table 6: Comparison with prior work. Our 23–32% success rate matches ML-DSA’s expected $\approx$20–25%. Time is average to successful signature (ML-DSA-44, $n=5$, $t=3$).
@@
Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility. The observed 23–32% success rate matches single-signer ML-DSA’s expected $\approx 20$$25\%$, which is inherent to the FIPS 204 specification.
@@
Table 7 quantifies the overhead of threshold signing compared to standard (single-party) ML-DSA across all three security levels.
Security Level Config Std Sign Thresh Sign Overhead Success
@@
5.7 Large-Scale Evaluation
We validated scalability with $n\in\{50,75,100\}$ parties (ML-DSA-44, P1-TEE):
  • $n=50$: 3.6s mean, 2.7 attempts
  • $n=75$: 6.1s mean, 2.6 attempts
  • $n=100$: 7.9s mean, 2.1 attempts
Signing scales linearly at $\approx$75–80 ms/party with consistent rejection rate.
@@ 5.8 Complexity Analysis
Per-party: $O(n\cdot(\ell+k))$ for shares, $O(N)$ for seeds.
@@ 5.9 FIPS 204 Compliance Verification
We verified our implementation against FIPS 204 specifications:
  • Parameter verification: All 19 FIPS 204 constants validated ($q=8380417$, $n=256$, $d=13$, etc.)
  • Algorithm correctness: 117/117 test cases passed covering KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling)
  • Bound verification: All signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$
NIST ACVP test vectors are publicly available; full byte-level KAT compliance requires implementing the FIPS 204 Section 7.2 encoding functions, which we leave as future work for production deployment.
@@ 5.10 Code Availability
A reference implementation ($\approx$2500 lines Python) with test suite and benchmarks will be released upon publication.
@@ 6.2 Optimistic Blame Protocol
  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.8)^{30}<0.13\%$)
@@ 6.3 Profile P2: Fully Distributed Signing
Our protocol achieves 8 online rounds:
Rounds Operation
1–2 Share exchange and aggregation
3–4 Masked reveal (commit-then-open)
@@ 5 A2B conversion via edaBits
6–7 Binary comparison
8 AND aggregation and output
@@
  • Our P2: 8 online rounds, dishonest majority, arbitrary $T$
@@ 6.4 Security Analysis of Profile P2
6.5 Profile P3: 2PC-Assisted Signing
Profile P3 designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC. Each CP receives additive shares: $\mathbf{S}_{1}+\mathbf{S}_{2}=c\mathbf{s}_{2}$. Using garbled circuits [Yao86], they evaluate the comparison in 2–3 rounds without reconstructing $c\mathbf{s}_{2}$.
Protocol Structure.
@@ 6.5 Profile P3: 2PC-Assisted Signing
6.5.1 UC Security Analysis of Profile P3
We prove that P3 achieves UC security under the assumption that at least one CP is honest.
@@ 6.5.1 UC Security Analysis of Profile P3
Theorem 6.5 (Profile P3 UC Security).
@@ Conclusion.
Corollary 6.7 (Full P3 Security).
Profile P3 achieves:
@@
[…]

### v3

@@ 4.7.3 Summary of UC Security
  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$ Achieves 3 rounds.
  • Profile P2 (MPC): UC security against dishonest majority via MPC (Theorem 6.2). Achieves 5 online rounds through combiner-mediated commit-then-open and constant-depth comparison circuits. This provides the strongest cryptographic guarantees with no hardware trust.
  • Profile P3+ (2PC Semi-Async): UC security under 1-of-2 CP honest assumption (Theorem 6.5). Achieves 2 logical rounds (1 signer step + 1 server round) with semi-asynchronous signer participation. Signers precompute nonces offline and respond within a time window. This profile is designed for human-in-the-loop authorization scenarios.
The choice of profile affects trust assumptions and performance, but not the fundamental cryptographic security of the resulting signatures. Table 4 summarizes the trade-offs.
Property P1 (TEE) P2 (MPC) P3+ (2PC)
Online rounds 3 5 2
Trust assumption TEE ideal None 1/2 CP honest
Dishonest majority Via TEE Via 2PC
Semi-async signers $\times$ $\times$
UC security
Table 4: Security profile comparison. P3+ = 1 signer step + 1 server round.
@@ 5.1 Implementation Overview
The implementation consists of several core modules. The polynomial arithmetic module provides operations over $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$ with $q=8380417$, including addition, subtraction, multiplication, and NTT-based fast multiplication. The Shamir secret sharing module handles polynomial evaluation and Lagrange interpolation over $\mathbb{Z}_{q}$; Lagrange coefficients $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$ depend on the signing set $S$ and are computed per-session in $O(T^{2})$ time, though they can be precomputed and cached for fixed-$S$ deployments. The mask generation module computes pairwise-canceling masks using SHAKE-256 as the PRF with efficient seed management. The protocol logic module implements profile-dependent signing with commitment verification, challenge derivation, and signature aggregation. Profile-specific modules include the preprocessing module (P2) for EdaBit and Beaver triple generation, the pre-signing module (P3+) for offline nonce precomputation, and the GC precomputation module (P3+) for offline garbled circuit generation.
@@
[…]
@@ 5.2 Benchmark Methodology
All benchmarks were conducted on an Intel Core i7-12700H (14 cores, 20 threads) with 32 GB DDR5-4800 RAM, running Ubuntu 22.04 LTS. Our Rust 1.75+ implementation uses Rayon for parallelization.
Our Rust implementation ($\approx$4000 lines) provides optimized performance through parallel polynomial operations using Rayon, NTT-based polynomial multiplication with Montgomery reduction, and efficient EdaBit and Beaver triple pooling with lazy generation.
@@
We report averages over 5 independent runs, each consisting of 5 trials per configuration (25 total measurements per data point).
@@ 5.3 Performance Results
Table 5 shows signing performance for various threshold configurations across all three protocol profiles, measured with our optimized Rust implementation.
P1 (TEE) P2 (MPC) P3+ (2PC)
$T$ NN Time Att. Time Att. Time Att.
2 3 3.9 ms 3.2 46.3 ms 4.1 16.7 ms 2.7
3 5 5.8 ms 3.2 21.5 ms 3.0 22.1 ms 3.7
5 7 10.3 ms 3.4 73.4 ms 4.4 24.0 ms 3.5
7 11 12.0 ms 2.9 20.5 ms 2.2 29.4 ms 3.4
11 15 30.3 ms 4.7 73.9 ms 4.3 39.2 ms 3.6
16 21 31.3 ms 3.4 194.4 ms 4.8 49.6 ms 2.8
21 31 33.1 ms 2.7 149.3 ms 4.0 62.2 ms 3.1
32 45 60.2 ms 3.2 129.3 ms 2.7 93.5 ms 3.6
Table 5: Signing performance across threshold configurations (ML-DSA-65, Rust implementation, Intel i7-12700H). Values are averages over 5 independent runs, each with 5 trials per configuration. Due to rejection sampling variance, P2 exhibits high variance (up to $14\times$ between runs).
Scaling Analysis.
P1 (TEE) scales from 4 ms at $T=2$ to 60 ms at $T=32$ ($\approx 15\times$), dominated by polynomial operations linear in $T$. P2 (MPC) shows higher variance (21–194 ms) because total time equals per-attempt time times retry count; notably, P2 latency for $T=32$ (129 ms) is lower than $T=16$ (194 ms) purely due to the probabilistic nature of rejection sampling—the $T=32$ runs averaged only 2.7 attempts versus 4.8 for $T=16$, while per-attempt latency scales as expected. P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency; the 2PC cost is largely independent of threshold.
@@ 5.4 Protocol Profile Comparison
Table 6 compares the three deployment profiles with experimental measurements for ML-DSA-65, $(n=5,t=3)$ using our optimized Rust implementation.
Profile Rounds Trust Async Online Time Attempts
P1-TEE 3 TEE $\times$ 5.8 ms 3.2
P2-MPC 5 $t$-of-$n$ $\times$ 21.5 ms 3.0
P3+ 2 1-of-2 CP 22.1 ms 3.7
Table 6: Protocol profile comparison (ML-DSA-65, $n=5$, $t=3$, Rust implementation). P3+ = 1 signer step + 1 server round; signers precompute nonces offline.
Analysis.
Profile P1 (TEE) is fastest at 5.8 ms but requires a trusted execution environment, making it suitable for HSM-based deployments. Profile P2 (5-round MPC) achieves 21.5 ms online time for 3-of-5 through parallel EdaBit-based A2B conversion and constant-depth AND tree for comparison; the 5-round structure merges commit-then-open (Rounds 3–4 $\to$ Round 2) and parallelizes comparison (Rounds 6–7 $\to$ Round 4). Profile P3+ provides 22.1 ms latency with FROST-like workflow: signers precompute nonces offline (2.5 KB per set) and respond within a time window (default: 5 min), requiring only a single user interaction. With GC precomputation, online 2PC evaluation drops to $\approx$12 ms.
@@ 5.5 Comparison with Prior Work
Scheme $T$ Size Rnd FIPS Async Time
dPN25 [dPN25] 8 3.3 KB 3 Yes $\times$
Noise flood [BCDG21] $\infty$ 17 KB 3 No $\times$ 300 ms
Ringtail [BMMSZ25] $\infty$ 13 KB 2 No $\times$ 200 ms
Ours (P1) $\infty$ 3.3 KB 3 Yes $\times$ 5.8 ms
Ours (P2) $\infty$ 3.3 KB 5 Yes $\times$ 21.5 ms
Ours (P3+) $\infty$ 3.3 KB 2 Yes 22.1 ms
Table 7: Comparison with prior work. Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, $n=5$, $t=3$). Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%. P3+ = 1 signer step + 1 server round.
@@
Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility. The observed $\approx$21–45% success rate is comparable to single-signer ML-DSA’s expected $\approx$20–25%, inherent to the FIPS 204 specification.
@@
Table 8 quantifies the overhead of threshold signing compared to standard (single-party) ML-DSA for ML-DSA-65.
Config Profile Std Sign Thresh Sign Overhead Attempts
@@
Table 8: Threshold signing overhead vs. standard ML-DSA-65 (Rust implementation). Standard ML-DSA signing takes $\approx$0.5 ms in optimized Rust. Overhead scales with $T$ but remains practical even at $T=32$.
@@
5.7 Scalability Analysis
Our benchmarks demonstrate practical scalability up to $T=32$ signers. Profile P1-TEE scales from 4 ms ($T=2$) to 60 ms ($T=32$), approximately linear in $T$. Profile P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency throughout. Profile P2-MPC shows higher variance (21–194 ms) due to MPC overhead sensitivity to retry count, exceeding 100ms for $T\geq 16$.
For larger thresholds ($T>32$), the dominant cost becomes polynomial operations ($O(T\cdot n\cdot\log n)$ per signature). Extrapolating from our measurements, $T=64$ would require approximately 100–200 ms for P1-TEE and 150–250 ms for P3+.
5.8 Preprocessing Analysis
Profile P2 (MPC) benefits significantly from preprocessing. Table 9 shows the amortization effect.
Component Per-Attempt Offline Gen. Storage
EdaBits 1536 45 ms 6.1 KB
Beaver triples $\sim$50,000 180 ms 200 KB
Total (1 set) 225 ms 206 KB
Buffer (5 sets) 1.13 s 1.03 MB
Table 9: P2 preprocessing requirements (ML-DSA-65, 5 parties, Rust implementation). Online consumption is <<1 ms per set after preprocessing.
P2 Implementation Optimizations.
Our Rust implementation achieves 21.5 ms online time (3-of-5) through several optimizations. Using Rayon, all 1536 coefficient A2B conversions run in parallel across CPU cores, and each level of the comparison AND tree executes pairs in parallel. EdaBits are only generated when the pool falls below threshold (conditional preprocessing), avoiding redundant computation. Beaver triples are pre-fetched per AND-tree level for cache efficiency.
The main bottleneck remains MPC communication overhead (simulated as computation in our single-machine benchmark). In a distributed deployment, network latency would dominate, making preprocessing even more critical.
5.9 Pre-signing Analysis (P3+)
Profile P3+ enables semi-asynchronous signing through offline nonce precomputation.
Component Size Generation Time
Nonce $\mathbf{y}_{i}$ 1.5 KB 8 ms
Commitment $\mathbf{w}_{i}$ 768 B 12 ms
Commitment hash 32 B <<1 ms
Total (1 set) 2.3 KB 21 ms
Buffer (10 sets) 23 KB 210 ms
Table 10: P3+ pre-signing storage and generation (ML-DSA-65, per party). Precomputed sets are consumed on signing; buffer replenishment runs in background.
GC Precomputation.
The r0-check garbled circuit has $\approx$1.2M gates. Pre-garbling takes $\approx$45 ms; online evaluation takes $\approx$12 ms. With precomputation, the 2PC phase drops from 80 ms to 15 ms.
5.10 Retry Efficiency Analysis
Due to rejection sampling, signing requires multiple attempts. Table 11 analyzes retry behavior across profiles (3-of-5 configuration).
Profile $p_{\text{succ}}$ $\mathbb{E}[\text{att}]$ Per-Att Time Retry Cost PP(5 att)
P1-TEE 31% 3.2 1.8 ms Full 85%
P2-MPC 33% 3.0 7.2 ms Full 87%
P3+ 27% 3.7 6.0 ms Partial 79%
Table 11: Retry efficiency (3-of-5, ML-DSA-65). $p_{\text{succ}}$ = per-attempt success rate (inherent to FIPS 204 rejection sampling); $\mathbb{E}[\text{att}]$ = expected attempts; $P(5)$ = success probability within 5 attempts. P3+ retries reuse precomputed nonces, reducing per-retry latency.
Profile P3+ has a unique advantage: failed attempts only consume the 2PC evaluation cost ($\approx$10 ms with precomputation), while nonce regeneration happens asynchronously in the background. This makes P3+ particularly efficient for high-throughput scenarios where multiple signatures are needed in succession.
@@ 5.11 Complexity Analysis
Per-party: $O(n\cdot(\ell+k))$ for shares, $O(N)$ for seeds. P3+ adds $O(B\cdot(\ell+k))$ for pre-signing buffer of size $B$.
@@ 5.12 FIPS 204 Compliance Verification
We verified our implementation against FIPS 204 specifications. Parameter verification confirmed all 19 FIPS 204 constants ($q=8380417$, $n=256$, $d=13$, etc.). Algorithm correctness testing covered KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling) with 117 test cases. Bound verification confirmed all signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$.
Our “FIPS 204-Compatible” claim refers to signature format compatibility: threshold-generated signatures are byte-identical to single-signer ML-DSA signatures and pass any compliant FIPS 204 verifier without modification. Full NIST ACVP Known-Answer-Test (KAT) certification for production deployment would additionally require implementing the deterministic encoding functions specified in FIPS 204 Section 7.2.
@@ 5.13 Code Availability
We provide two implementations:
  • Python reference ($\approx$3500 lines): Readable implementation for protocol understanding and verification.
[…]
Both implementations include:
  • All three protocol profiles (P1-TEE, P2-MPC 5-round, P3+-SemiAsync)
  • Preprocessing modules for P2 (EdaBits, Beaver triples with parallel generation)
  • Pre-signing and GC precomputation modules for P3+
  • Comprehensive protocol test suite with 40+ integration tests (separate from 117 FIPS compliance tests)
  • Benchmark harness for reproducible measurements across threshold configurations
Code will be released upon publication.
@@ 6.2 Optimistic Blame Protocol
  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.75)^{30}<10^{-4}$)
@@ 6.3 Profile P2: Fully Distributed Signing
Using combiner-mediated commit-then-open and constant-depth comparison circuits, our optimized protocol achieves 5 online rounds:
Round Operation
1 Share exchange + aggregation
2 Combiner-mediated masked reveal
@@ 3 A2B conversion via edaBits
4 Parallel comparison + AND aggregation
5 Output
Round Optimization Techniques.
  1. 1.
    Combiner-mediated commit-then-open (Rounds 3-4 $\to$ Round 2): All parties send $(Com_{i},masked_{i},rand_{i})$ to a designated combiner, who releases all openings simultaneously after collecting from all parties. SPDZ MAC check catches any tampering by the combiner.
  2. 2.
    Constant-depth comparison (Rounds 6-7 $\to$ Round 4): All 1536 coefficient comparisons are evaluated in parallel (Layer 1), followed by an AND-tree aggregation (Layer 2). Using Beaver triples, all $(d,e)$ values can be opened in a single round.
@@ 5 Output
  • Our P2: 5 online rounds, dishonest majority, arbitrary $T$
@@ 6.4 Security Analysis of Profile P2
6.5 Profile P3+: Semi-Async 2PC-Assisted Signing
Profile P3+ designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC. The key innovation is semi-asynchronous signer participation: signers precompute nonces offline and respond within a time window, rather than participating in multiple synchronous rounds.
Semi-Async Timing Model.
  1. 1.
    Pre-signing Window $t_{0},t_{1}]$: Fully asynchronous. Each signer precomputes $B$ sets of $(\mathbf{y}_{i},\mathbf{w}_{i},\mathsf{Com}_{i})$ at any time and broadcasts commitments to the combiner.
  2. 2.
    Challenge Broadcast $t_{2}$: Combiner aggregates commitments, derives challenge $c$, and broadcasts to signers.
  3. 3.
    Response Window $t_{2},t_{2}+\Delta]$: Semi-synchronous. Signers respond within $\Delta$ (e.g., 5 minutes). No coordination between signers required.
  4. 4.
    2PC Execution $t_{3},t_{4}]$: Synchronous between CPs only.
Signer Workflow.
Upon receiving a signing request, a signer’s client daemon retrieves a precomputed nonce set and computes the response automatically. Retries due to rejection sampling are handled transparently by the daemon. The entire process completes within $\sim$1 second of authorization, requiring only a single user interaction.
2PC Protocol Structure.
@@ 6.5 Profile P3+: Semi-Async 2PC-Assisted Signing
GC Precomputation.
Since the r0-check circuit structure is fixed (independent of inputs), $\mathsf{CP}_{1}$ can pre-garble circuits offline. Online, only OT and evaluation are needed, reducing 2PC latency from $\sim$50ms to $\sim$10–20ms.
6.5.1 UC Security Analysis of Profile P3+
We prove that P3+ achieves UC security under the assumption that at least one CP is honest.
@@ 6.5.1 UC Security Analysis of Profile P3+
Theorem 6.5 (Profile P3+ UC Security).
@@ Conclusion.
Corollary 6.7 (Full P3+ Security).
Profile P3+ achieves:
@@
P3+ achieves 22 ms latency for 3-of-5 threshold (comparable to P2’s 22 ms, but with semi-async signer participation). The 2PC overhead is minimal compared to P2’s MPC because only 2 parties participate in the 2PC subprotocol, garbled circuits have constant rounds, and no edaBits preprocessing is required. With GC precomputation, 2PC latency drops to $\sim$12 ms.
Round Complexity Analysis.
Profile Rounds/Attempt Expected Total
P2 (5-round) 5 15
P3+ 2 7.4
P3+ logical rounds = 1 signer step + 1 server round. Expected total assumes $\approx$30% per-attempt success rate (3.7 attempts for P3+, 3.0 for P2).
Comparison with Concurrent Work.
The “Async” column in our comparison table is unique to P3+. No concurrent work supports semi-asynchronous signer participation; signers in all other schemes must synchronize for multiple rounds. P3+ is suited for human-in-the-loop authorization workflows such as multi-party transaction authorization and distributed key custody.